Find the partial differential equation arising from f[z/x^3,y/z]=0, where f is an arbitrary function of the arguments. Also find the general solution of the PDE obtained.
1
Expert's answer
2020-03-18T18:21:06-0400

Let "F(x,y,z)=f\left(\frac{z}{x^3},\frac{y}{z}\right)". Then "F'_x=f'_{x_1}\left(-3\frac{z}{x^4}\right)+f'_{x_2}\cdot 0=-3f'_{x_1}\frac{z}{x^4}", "F'_y=f'_{x_1}\cdot 0+f'_{x_2}\frac{1}{z}=f'_{x_2}\frac{1}{z}", "F'_z=f'_{x_1}\frac{1}{x^3}-f'_{x_2}\frac{y}{z^2}".

Since our function "z(x,y)" is given by the equation "F(x,y,z)=0", by the implicit function theorem we obtain "z'_x=-\frac{F'_x}{F'_z}=\frac{3f'_{x_1}\frac{z}{x^4}}{f'_{x_1}\frac{1}{x^3}-f'_{x_2}\frac{y}{z^2}}" and "z'_y=-\frac{F'_y}{F'_z}=\frac{-f'_{x_2}\frac{1}{z}}{f'_{x_1}\frac{1}{x^3}-f'_{x_2}\frac{y}{z^2}}".

Rewrite it: "\frac{x}{3z}z'_x=\frac{f'_{x_1}\frac{1}{x^3}}{f'_{x_1}\frac{1}{x^3}-f'_{x_2}\frac{y}{z^2}}" and "\frac{y}{z}z'_y=\frac{-f'_{x_2}\frac{y}{z^2}}{f'_{x_1}\frac{1}{x^3}-f'_{x_2}\frac{y}{z^2}}", so "\frac{x}{3z}z'_x+\frac{y}{z}z'_y=1" or "xz'_x+3yz'_y=3z".


For obtain a general solution of this equation, we write the following equation: "\frac{dx}{x}=\frac{dy}{3y}=\frac{dz}{3z}".

1)Solve "\frac{dx}{x}=\frac{dz}{3z}" . We have "3zdx-xdz=0". Multiplying it by "\frac{x^2}{z^2}" , we obtain "\frac{3x^2zdx-x^3dz}{z^2}=0" or "d\left(\frac{x^3}{z}\right)=0". So we obtain a first integral "\frac{x^3}{z}=C_1".

2)Solve "\frac{dy}{3y}=\frac{dz}{3z}" . We have "ydz-zdy=0". Multiplying it by "\frac{1}{z^2}", we obtain "\frac{ydz-zdy}{z^2}=0" or "-d\left(\frac{y}{z}\right)=0" . So we obain a second integral "\frac{y}{z}=C_2".


Therefore the general solution of "xz'_x+3yz'_y=3z" is "g\left(\frac{x^3}{z},\frac{y}{z}\right)=0"